# What Language Models Know But Don’t Say: Non-Generative Prior Extraction for Generalization

### v1

@@ A.5.2 Computational Efficiency
where $N_{\text{rounds}}\in[3,5]$ and $N_{\text{queries}}\in[5,10]$. This results in more than 28 LLM queries per feature on average.
@@
A.6 Prior Impact on Model Performance
A.6.1 LoID Priors vs. $\mathcal{N}$(0, 0.45) vs. $\mathcal{N}$(0, 1)
To isolate the contribution of the prior means inferred by LoID, we evaluate a baseline using normal priors with zero mean and standard deviation $0.45$ (the Goldilock setting). This comparison quantifies how much LoID’s learned prior means contribute to downstream AUC. As shown in Table 11, incorporating LoID-derived means yields a substantial performance gain, increasing the gap closed by approximately 154% on average. This difference is most prominent in datasets where the uninformed priors catastrophically fail: Occupancy (+1156%), Diabetes (+105%), and Pima (+98%). In these cases, the $\mathcal{N}(0,0.45)$ prior provides insufficient directional guidance, causing the Bayesian posterior to converge toward coefficients near zero and perform worse than the OOD baseline. Conversely, LoID’s informative means—despite systematic positive bias—capture feature importance magnitudes that guide the model toward the correct solution space. Even on datasets where LoID underperforms (Heart: -32%, Liver: -10%), the performance gap remains small compared to the massive gains elsewhere. This demonstrates that prior mean quality matters substantially more than perfect directional correctness when paired with appropriately calibrated uncertainty ($\sigma\approx 0.4$$0.5$).
Finally, we include an additional $\mathcal{N}(0,1)$ baseline to illustrate the improvement achieved over a more weakly informative prior that is commonly used with Bayesian inference.
A.6.2 $\mathcal{N}$(0, 1) vs. $\mathcal{U}$(-1, 1)
Table 12 shows the catastrophic failure of Uniform priors (average AUC 0.42, with Occupancy at 0.02) and reveals that while Uniform priors ostensibly provide “maximum uninformativeness” by assigning equal probability across a range, they fundamentally fail as regularizers for regression coefficients because:
  1. 1.
  2. 2.
    Incompatibility with MCMC: The flat density of Uniform priors provides no gradient information to guide sampling algorithms, resulting in inefficient exploration and convergence issues, particularly visible in the Occupancy dataset collapse.
  3. 3.
    Prior-likelihood mismatch: For logistic regression, coefficients naturally follow a scale where values near zero are common and extreme values are rare. Uniform priors violate this structure, treating $\beta=0.1$ and $\beta=4.9$ as equally plausible a prior.
Dataset $\mathcal{N}(0,1)$ $\mathcal{U}(-1,1)$
Bank 0.61 0.35
Blood 0.63 0.59
Credit Score 0.74 0.58
Diabetes 0.82 0.32
Heart 0.91 0.84
Income 0.81 0.30
Liver 0.68 0.70
Occupancy 0.72 0.02
Jungle 0.72 0.58
Pima 0.82 0.32
Average 0.75 0.46
Table 12: Comparison of LoID priors against uninformative Normal and Uniform priors on OOD datasets.
A.7 Computational Resources

### v2

@@ A.7 Computational efficiency
where $N_{\text{rounds}}=10$ and $N_{\text{queries}}=10$. This results in more than 28 LLM queries per feature on average.
@@
A.8 Resources and reproducibility
Code availability
Data availability
All datasets used in this study except Head and Neck Cancer Recurrence are publicly available through open-access repositories. Further details and references are provided in Section A.4. All preprocessing steps required to obtain the final train/test splits used in our experiments, along with the preprocessed data, are available in our code release.
The Head and Neck Cancer Recurrence dataset will be released upon publication.
Results variability
Computational resources